OpenStudy (anonymous):

Which of the following equations is the result of completing the square on x2 - 6x - 9 = 0

OpenStudy (anonymous):

x^2-6x-9=0 , transfer -9 to the right .. then divide the value of b and square it after; where b=-6 . it will be like this: x^2-6x+__=0+9+__

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

could you draw it out because i dont understand it that way

OpenStudy (anonymous):

okay .. BRB

OpenStudy (anonymous):